OpenStudy (anonymous):

what is the intercetion point of the graphs of the equations f(x) = 4x -1 and g(x) = x + 2

OpenStudy (anonymous):

4x-1=x+2 3x=3 x=1 y = x+2 = 3 hence the pt is (1,3)